# Sub-GeV Dark Matter Direct Detection with Neutrino Observatories

### v1

@@ Abstract
Beyond JUNO, our idea can be applied to other neutrino observatories, such as Super-K/Hyper-K, DUNE, and IceCube. While these detectors have less light emission than JUNO, they also have lower dark rates. Utilizing multiple detectors would serve two key purposes. First, making searches more robust by comparing different detectors, as they should observe consistent amplitudes and phases for an annual modulation signal, which does not depend on detector location, though some backgrounds do. Second, extracting more detailed DM properties. If a signal were observed in JUNO, we would extract only a degenerate combination of DM mass and cross section, which would appear as a characteristic trajectory in Fig. 2. Independent observations could help disentangle the specific DM mass and cross section, as the trajectories depend on the detector properties. It would also be interesting to explore the potential of this technique for dedicated DM detectors. We will explore these topics in future work.
@@
Ultimately, this work opens a new frontier in DM detection. Neutrino observatories, which are designed for completely different physics, may hold the key to uncovering sub-GeV DM. The potential is now clear, but realizing it will require dedicated experimental validation, sharper theoretical modeling, and coordinated multi-detector searches. If DM hides in this mass range, we may already have the tools to find it. We just need to start looking, but in a new way.
@@
Acknowledgments.— We are grateful for helpful discussions with Carlos Blanco, Yufeng Li, Bernhard Mistlberger, Mark Vagins, and Tien-Tien Yu. RKL was supported by the U.S. Department of Energy under Contract DE-AC02-76SF00515. JFB was supported by National Science Foundation Grant No. PHY-2310018.

### v2

@@ Abstract
Beyond JUNO, this new technique can be generalized to other neutrino observatories, both scintillator-based (those noted above plus Daya Bay, KamLAND, and more) and otherwise (Super-K/Hyper-K, IceCube, and DUNE). It would also be interesting to explore the potential of this technique for dedicated DM detectors. Utilizing multiple detectors is important because they will have very different backgrounds and systematics. Further, if a signal were observed in JUNO, they would extract only a degenerate combination of DM mass and cross section — a characteristic trajectory in Fig. 2, whereas other detectors would have different trajectories, which could cross JUNO’s. In future work, we will address these and other topics.
@@
Ultimately, this work opens a new frontier in DM detection. Neutrino observatories, which are designed for completely different physics goals, may hold the key to uncovering sub-GeV DM. The potential is now clear, but realizing it will require dedicated experimental validation, sharper theoretical modeling, and coordinated multi-detector searches. If DM hides in this mass range, we may already have the tools to find it. We just need to start looking, but in a new way.
@@
Acknowledgments.— We are grateful for helpful discussions with Carlos Blanco, Rouven Essig, Roni Harnik, Noah Kurinsky, Yufeng Li, Bryce Littlejohn, Stephan Meighen-Berger, Bernhard Mistlberger, Mark Vagins, and Tien-Tien Yu. RKL was supported by the U.S. Department of Energy under Contract DE-AC02-76SF00515. JFB was supported by National Science Foundation Grant No. PHY-2310018.